OpenStudy (anonymous):

*Can anyone explain how to do this?* A cyclist rides her bike at a rate of 12 meters per second. What is this rate in kilometers per hour? How many kilometers will the cyclist travel in 4 hours?

OpenStudy (anonymous):

You need to first find out how much seconds are in a minute so 60. times 12x60 and you will have one minute. Times that number by 60 to get one hour. Time that one hour by 4 to get you total of hoe much she rides in 4 hours.

OpenStudy (anonymous):

*how

OpenStudy (anonymous):

Do you need the answer

OpenStudy (anonymous):

That would be meters per hour

OpenStudy (anonymous):

Oh well, just convert those meters in kilometers and thats your answer

OpenStudy (anonymous):

For every 1000m, it is 1 kilometer

OpenStudy (anonymous):

Example, 1200m, would be 1.2 kilometers

OpenStudy (anonymous):

:O, what happened to my medal?!?

OpenStudy (anonymous):

I took it back because palmer is about to unload an information bomb on us all

OpenStudy (anonymous):

I'm waiting for whpalmer's response. I would like to know what it is.

OpenStudy (whpalmer4):

12 m/s * s/m * m/hr \[1 \text{ minute} = 60 \text{ seconds}\]Divide both sides by 1 minute: \[1 = \frac{60 \text{ seconds}}{1 \text{ minute}}\] \[1 \text{ hour} = 60 \text{ minutes}\]Divide both sides by 1 hour: \[1 = \frac{60 \text{ minutes}}{1 \text{ hour}}\] \[1 \text{ kilometer} = 1000 \text{ meters}\]Divide both sides by 1000 meters: \[1 = \frac{1 \text{ kilometer}}{1000 \text{ meters}}\] You can multiply anything by 1 without changing its value, so we'll multiply \[12 \frac{\text{ meters}}{\text{ second}} * \frac{60 \text{ minutes}}{1 \text{ hour}} * \frac{ 60 \text{ seconds}} {1 \text{ minute}}\frac{1 \text{ kilometer}}{1000 \text{ meters}} = \]\[12 \frac{\cancel{\text{ meters}}}{\cancel{\text{ second}}} * \frac{60 \cancel{\text{ minutes}}}{1 \text{ hour}} * \frac{ 60 \cancel{\text{ seconds}}} {1 \cancel{\text{ minute}}}* \frac{1 \text{ kilometer}}{1000 \cancel{\text{ meters}}}\]\[= \text{_______ kilometers/hour}\]

OpenStudy (whpalmer4):

If you do your conversions like this, you get built-in error checking. If the units don't cancel properly, you've set up the conversion wrong.

OpenStudy (whpalmer4):

It's a bit more writing, but so is having to do the problem over because you muffed the unit conversion :-)

OpenStudy (anonymous):

Oh, wow. Thank you (: I had originally thought the question was missing information, but you certainly helped whpalmer.

OpenStudy (whpalmer4):

Great! You're the lucky winner of my last helping effort for the evening, time to go to rehearsal :-)